# Danny Halawi4, Fred Zhang4, Philip E. Tetlock2,5

### v2

@@ Appendix H Top 50 ForecastBench Leaderboards
Superforecaster median forecast ForecastBench 0.097 0.04 0.068 [0.06, 0.077] 0%
Public median forecast ForecastBench 0.127 0.039 0.083 [0.075, 0.091] $<$0.001 28%
GPT-4 OpenAI Freeze values Zero shot 0.131 0.041 0.086 [0.077, 0.095] $<$0.001 34%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.135 0.04 0.088 [0.079, 0.097] $<$0.001 33%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.114 0.062 0.088 [0.076, 0.1] $<$0.001 34%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.118 0.064 0.091 [0.08, 0.102] $<$0.001 28%
GPT-4o OpenAI Freeze values Scratchpad 0.134 0.052 0.093 [0.084, 0.102] $<$0.001 28%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.13 0.06 0.095 [0.085, 0.105] $<$0.001 28%
GPT-4o OpenAI News with freeze values Scratchpad 0.142 0.049 0.096 [0.087, 0.104] $<$0.001 27%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.134 0.062 0.098 [0.088, 0.108] $<$0.001 25%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.13 0.068 0.099 [0.091, 0.107] $<$0.001 24%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.146 0.057 0.102 [0.091, 0.112] $<$0.001 30%
Claude-3-5-Sonnet-20240620 Anthropic No Scratchpad 0.118 0.086 0.102 [0.09, 0.113] $<$0.001 26%
Claude-2.1 Anthropic Freeze values Scratchpad 0.163 0.042 0.102 [0.095, 0.11] $<$0.001 29%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.139 0.067 0.103 [0.093, 0.114] $<$0.001 26%
GPT-4o OpenAI No Scratchpad 0.134 0.073 0.104 [0.094, 0.113] $<$0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News No 0.135 0.074 0.104 [0.094, 0.114] $<$0.001 25%
Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.153 0.056 0.105 [0.096, 0.114] $<$0.001 26%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.135 0.076 0.105 [0.096, 0.115] $<$0.001 24%
GPT-4 OpenAI Freeze values Scratchpad 0.151 0.061 0.106 [0.097, 0.114] $<$0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.13 0.083 0.107 [0.096, 0.118] $<$0.001 24%
Claude-2.1 Anthropic No Scratchpad 0.163 0.052 0.107 [0.099, 0.116] $<$0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic No Zero shot 0.114 0.101 0.107 [0.094, 0.12] $<$0.001 26%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.151 0.064 0.108 [0.099, 0.116] $<$0.001 23%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.154 0.066 0.11 [0.099, 0.121] $<$0.001 29%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.151 0.069 0.11 [0.1, 0.12] $<$0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic News No 0.136 0.085 0.11 [0.099, 0.121] $<$0.001 24%
GPT-4o OpenAI Freeze values Zero shot 0.166 0.055 0.11 [0.1, 0.121] $<$0.001 30%
GPT-4o OpenAI News Scratchpad 0.142 0.08 0.111 [0.1, 0.122] $<$0.001 23%
Gemini-1.5-Pro Google Freeze values Zero shot 0.154 0.068 0.111 [0.099, 0.123] $<$0.001 28%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.139 0.084 0.111 [0.102, 0.121] $<$0.001 23%
Claude-3-Opus-20240229 Anthropic No Scratchpad 0.135 0.088 0.111 [0.102, 0.121] $<$0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.152 0.072 0.112 [0.101, 0.122] $<$0.001 26%
GPT-4-Turbo-2024-04-09 OpenAI No Zero shot 0.135 0.089 0.112 [0.102, 0.123] $<$0.001 22%
GPT-4 OpenAI No Zero shot 0.131 0.095 0.113 [0.103, 0.122] $<$0.001 24%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.143 0.084 0.113 [0.103, 0.123] $<$0.001 22%
Mistral-Large-Latest Mistral AI No Scratchpad 0.13 0.097 0.113 [0.104, 0.123] $<$0.001 23%
@@
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.15 0.077 0.114 [0.105, 0.123] $<$0.001 23%
Gemini-1.5-Flash Google Freeze values Zero shot 0.15 0.078 0.114 [0.101, 0.127] $<$0.001 27%
Gemini-1.5-Pro Google No Scratchpad 0.139 0.09 0.115 [0.105, 0.124] $<$0.001 22%
GPT-4 OpenAI No Scratchpad 0.151 0.079 0.115 [0.106, 0.123] $<$0.001 21%
Imputed Forecaster ForecastBench 0.203 0.029 0.116 [0.107, 0.124] $<$0.001 32%
Gemini-1.5-Pro Google News Scratchpad 0.143 0.091 0.117 [0.107, 0.127] $<$0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI No Scratchpad 0.152 0.082 0.117 [0.108, 0.127] $<$0.001 21%
Claude-2.1 Anthropic Freeze values Zero shot 0.179 0.055 0.117 [0.107, 0.127] $<$0.001 29%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.161 0.075 0.118 [0.108, 0.129] $<$0.001 26%
Claude-3-Opus-20240229 Anthropic No Zero shot 0.146 0.093 0.119 [0.108, 0.131] $<$0.001 24%
Claude-3-Opus-20240229 Anthropic News No 0.134 0.106 0.12 [0.11, 0.131] $<$0.001 23%
Llama-3-70b-Chat-Hf Meta No Zero shot 0.139 0.102 0.12 [0.11, 0.131] $<$0.001 23%
Gemini-1.5-Flash Google No Scratchpad 0.151 0.09 0.12 [0.11, 0.13] $<$0.001 22%
@@
Pairwise p-value comparing to No. 1 (bootstrapped): The p-value calculated by bootstrapping the differences in overall score between each model and the best forecaster (the group with rank 1) under the null hypothesis that there’s no difference.
Pct. more accurate than No. 1: The percent of questions where this forecaster had a better overall score than the best forecaster (with rank 1).

### v3

@@ Appendix H Leaderboards: Top 50
Superforecaster median forecast ForecastBench 0.126 0.059 0.092 [0.073, 0.112] 0%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.139 0.089 0.114 [0.092, 0.136] <0.01 31%
Public median forecast ForecastBench 0.156 0.072 0.114 [0.095, 0.133] <0.001 24%
@@
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.144 0.094 0.119 [0.098, 0.141] <0.001 30%
GPT-4 OpenAI Freeze values Zero shot 0.164 0.079 0.121 [0.100, 0.143] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.145 0.099 0.122 [0.095, 0.149] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.170 0.077 0.124 [0.102, 0.145] <0.001 33%
GPT-4o OpenAI News with freeze values Scratchpad 0.168 0.082 0.125 [0.107, 0.144] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.139 0.113 0.126 [0.104, 0.148] <0.001 29%
GPT-4o OpenAI Freeze values Scratchpad 0.163 0.093 0.128 [0.108, 0.148] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.144 0.116 0.130 [0.109, 0.151] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.159 0.103 0.131 [0.112, 0.150] <0.001 25%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.170 0.096 0.133 [0.110, 0.156] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 1 0.153 0.113 0.133 [0.112, 0.154] <0.001 26%
GPT-4o OpenAI Scratchpad 0.163 0.103 0.133 [0.116, 0.150] <0.001 25%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.163 0.105 0.134 [0.116, 0.152] <0.001 24%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.173 0.095 0.134 [0.110, 0.158] <0.001 26%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.161 0.109 0.135 [0.118, 0.152] <0.001 23%
GPT-4o OpenAI News Scratchpad 0.168 0.107 0.137 [0.117, 0.158] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.179 0.104 0.141 [0.119, 0.163] <0.001 27%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.145 0.138 0.141 [0.115, 0.168] <0.001 24%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.189 0.095 0.142 [0.126, 0.158] <0.001 25%
Gemini-1.5-Pro Google Scratchpad 0.168 0.116 0.142 [0.124, 0.160] <0.001 23%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.168 0.116 0.142 [0.124, 0.160] <0.001 23%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.172 0.113 0.142 [0.123, 0.161] <0.001 22%
GPT-4 OpenAI Freeze values Scratchpad 0.178 0.107 0.143 [0.124, 0.162] <0.001 21%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.181 0.105 0.143 [0.118, 0.168] <0.001 26%
Claude-3-Opus-20240229 Anthropic Scratchpad 0.163 0.124 0.144 [0.126, 0.161] <0.001 22%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.170 0.118 0.144 [0.125, 0.163] <0.001 22%
Claude-3-Opus-20240229 Anthropic News Superforecaster 1 0.161 0.128 0.144 [0.124, 0.165] <0.001 23%
Claude-2.1 Anthropic Freeze values Scratchpad 0.213 0.077 0.145 [0.128, 0.162] <0.001 26%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.179 0.111 0.145 [0.129, 0.161] <0.001 23%
Gemini-1.5-Flash Google Freeze values Zero shot 0.183 0.107 0.145 [0.118, 0.172] <0.001 24%
Gemini-1.5-Pro Google News Scratchpad 0.172 0.119 0.145 [0.127, 0.164] <0.001 22%
GPT-4o OpenAI Freeze values Zero shot 0.201 0.091 0.146 [0.121, 0.170] <0.001 28%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.183 0.108 0.146 [0.129, 0.163] <0.001 21%
Gemini-1.5-Pro Google Freeze values Zero shot 0.186 0.106 0.146 [0.120, 0.172] <0.001 25%
Qwen1.5-110B-Chat Qwen Freeze values Zero shot 0.195 0.097 0.146 [0.125, 0.167] <0.001 22%
GPT-4 OpenAI Scratchpad 0.178 0.116 0.147 [0.132, 0.162] <0.001 18%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.188 0.108 0.148 [0.126, 0.170] <0.001 26%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Zero shot 0.189 0.111 0.150 [0.124, 0.176] <0.001 26%
Mistral-Large-Latest Mistral AI Scratchpad 0.161 0.143 0.152 [0.134, 0.169] <0.001 22%
GPT-4 OpenAI Zero shot 0.164 0.141 0.152 [0.134, 0.170] <0.001 24%
Claude-2.1 Anthropic Scratchpad 0.213 0.092 0.152 [0.134, 0.171] <0.001 23%
Imputed Forecaster ForecastBench 0.250 0.056 0.153 [0.136, 0.170] <0.001 28%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.193 0.116 0.154 [0.137, 0.172] <0.001 20%
Claude-2.1 Anthropic Freeze values Zero shot 0.220 0.090 0.155 [0.133, 0.177] <0.001 27%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.185 0.126 0.155 [0.137, 0.174] <0.001 22%
@@
Mixtral-8x22B-Instruct-V0.1 Mistral AI News with freeze values Scratchpad 0.197 0.115 0.156 [0.138, 0.174] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.188 0.125 0.156 [0.136, 0.177] <0.001 24%
@@
Table 19: Leaderboard: LLM question set (Top 50)
Brier Score $\downarrow$
Model Organization
Information
provided
Prompt
Dataset
(N=4,107)
Market
(N=907)
Overall
(N=5,014)
Confidence
Interval
Pairwise
p-value
comparing
to No. 1
Pct. more
accurate
than No. 1
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.171 0.062 0.116 [0.111, 0.122] 0%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.171 0.065 0.118 [0.112, 0.123] 0.243 42%
GPT-4o OpenAI Freeze values Scratchpad 0.191 0.054 0.123 [0.117, 0.128] <0.001 43%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.162 0.089 0.125 [0.120, 0.131] <0.001 35%
GPT-4o OpenAI News with freeze values Scratchpad 0.192 0.066 0.129 [0.123, 0.134] <0.001 39%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.167 0.094 0.131 [0.125, 0.136] <0.001 34%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.190 0.076 0.133 [0.126, 0.139] <0.001 41%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.176 0.092 0.134 [0.128, 0.139] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.171 0.097 0.134 [0.129, 0.139] <0.001 31%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.187 0.083 0.135 [0.129, 0.141] <0.001 32%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.171 0.101 0.136 [0.130, 0.142] <0.001 10%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.195 0.077 0.136 [0.130, 0.143] <0.001 42%
Gemini-1.5-Pro Google Scratchpad 0.162 0.111 0.137 [0.131, 0.142] <0.001 32%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.206 0.068 0.137 [0.131, 0.144] <0.001 41%
Gemini-1.5-Pro Google News Scratchpad 0.167 0.109 0.138 [0.132, 0.144] <0.001 32%
GPT-4 OpenAI Freeze values Scratchpad 0.196 0.083 0.139 [0.133, 0.145] <0.001 36%
Imputed Forecaster ForecastBench 0.250 0.034 0.142 [0.138, 0.145] <0.001 47%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.210 0.074 0.142 [0.137, 0.148] <0.001 35%
GPT-4o OpenAI Scratchpad 0.191 0.094 0.143 [0.137, 0.148] <0.001 31%
Gemini-1.5-Pro Google Freeze values Zero shot 0.219 0.066 0.143 [0.137, 0.149] <0.001 39%
GPT-4o OpenAI News Scratchpad 0.192 0.095 0.143 [0.137, 0.149] <0.001 31%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.207 0.084 0.146 [0.139, 0.152] <0.001 33%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.187 0.104 0.146 [0.140, 0.151] <0.001 30%
LLM Crowd ForecastBench News 0.240 0.052 0.146 [0.143, 0.149] <0.001 38%
Qwen1.5-110B-Chat Qwen Scratchpad 0.176 0.116 0.146 [0.141, 0.151] <0.001 29%
GPT-4 OpenAI Freeze values Zero shot 0.222 0.071 0.146 [0.140, 0.152] <0.001 38%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.199 0.096 0.148 [0.142, 0.153] <0.001 26%
LLM Crowd ForecastBench News 0.241 0.055 0.148 [0.145, 0.152] <0.001 38%
LLM Crowd ForecastBench News 0.241 0.055 0.148 [0.145, 0.152] <0.001 37%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.204 0.093 0.149 [0.143, 0.155] <0.001 28%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.193 0.107 0.150 [0.143, 0.157] <0.001 31%
Gemini-1.5-Pro Google News Superforecaster 1 0.188 0.112 0.150 [0.144, 0.156] <0.001 32%
GPT-4 OpenAI Scratchpad 0.196 0.105 0.150 [0.145, 0.155] <0.001 28%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.193 0.109 0.151 [0.145, 0.157] <0.001 29%
Gemini-1.5-Pro Google Zero shot 0.219 0.083 0.151 [0.145, 0.158] <0.001 36%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.206 0.098 0.152 [0.146, 0.159] <0.001 31%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.210 0.095 0.152 [0.147, 0.158] <0.001 27%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.216 0.090 0.153 [0.148, 0.158] <0.001 25%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.208 0.098 0.153 [0.148, 0.158] <0.001 28%
Claude-3-Opus-20240229 Anthropic Zero shot 0.190 0.120 0.155 [0.148, 0.161] <0.001 34%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.207 0.103 0.155 [0.150, 0.161] <0.001 26%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.195 0.116 0.156 [0.149, 0.163] <0.001 34%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.207 0.104 0.156 [0.149, 0.163] <0.001 31%
Gemini-1.5-Flash Google Freeze values Zero shot 0.232 0.082 0.157 [0.150, 0.164] <0.001 40%
GPT-4o OpenAI News Superforecaster 3 0.211 0.104 0.157 [0.152, 0.163] <0.001 27%
Mixtral-8x22B-Instruct-V0.1 Mistral AI Freeze values Scratchpad 0.211 0.105 0.158 [0.152, 0.164] <0.001 30%
Gemini-1.5-Flash Google Scratchpad 0.193 0.125 0.159 [0.152, 0.165] <0.001 28%
Llama-3-70b-Chat-Hf Meta Zero shot 0.207 0.111 0.159 [0.153, 0.164] <0.001 28%
Claude-2.1 Anthropic Scratchpad 0.236 0.085 0.160 [0.155, 0.166] <0.001 38%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 1 0.211 0.113 0.162 [0.156, 0.168] <0.001 30%
Notes:
Table 20: Leaderboard: Human question set with LLM question set combination questions (Top 50)
Brier Score $\downarrow$
Model Organization
Information
provided
Prompt
Dataset
(N=1,312)
Market
(N=301)
Overall
(N=1,613)
Confidence
Interval
Pairwise
p-value
comparing
to No. 1
Pct. more
accurate
than No. 1
Superforecaster median forecast ForecastBench 0.093 0.048 0.071 [0.062, 0.079] 0%
Public median forecast ForecastBench 0.123 0.057 0.090 [0.082, 0.099] <0.001 24%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Scratchpad 0.158 0.090 0.124 [0.112, 0.136] <0.001 24%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Scratchpad 0.163 0.085 0.124 [0.113, 0.136] <0.001 24%
GPT-4o OpenAI Freeze values Scratchpad 0.182 0.069 0.125 [0.115, 0.136] <0.001 25%
GPT-4o OpenAI News with freeze values Scratchpad 0.175 0.084 0.129 [0.118, 0.140] <0.001 24%
Gemini-1.5-Pro Google Freeze values Scratchpad 0.155 0.111 0.133 [0.122, 0.144] <0.001 22%
Claude-3-5-Sonnet-20240620 Anthropic News with freeze values Scratchpad 0.163 0.109 0.136 [0.124, 0.148] <0.001 20%
Gemini-1.5-Pro Google News with freeze values Scratchpad 0.157 0.117 0.137 [0.126, 0.147] <0.001 21%
GPT-4-Turbo-2024-04-09 OpenAI Scratchpad 0.163 0.113 0.138 [0.128, 0.148] <0.001 18%
Claude-3-5-Sonnet-20240620 Anthropic Freeze values Zero shot 0.177 0.101 0.139 [0.126, 0.152] <0.001 23%
Gemini-1.5-Pro Google Scratchpad 0.155 0.125 0.140 [0.130, 0.150] <0.001 20%
Gemini-1.5-Pro Google News Scratchpad 0.157 0.124 0.141 [0.129, 0.152] <0.001 21%
Claude-3-5-Sonnet-20240620 Anthropic Scratchpad 0.158 0.123 0.141 [0.129, 0.152] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Zero shot 0.178 0.104 0.141 [0.128, 0.154] <0.001 23%
Qwen1.5-110B-Chat Qwen Freeze values Scratchpad 0.172 0.112 0.142 [0.131, 0.152] <0.001 17%
GPT-4o OpenAI Scratchpad 0.182 0.102 0.142 [0.132, 0.152] <0.001 20%
GPT-4o OpenAI News Scratchpad 0.175 0.109 0.142 [0.131, 0.153] <0.001 20%
GPT-4-Turbo-2024-04-09 OpenAI Freeze values Zero shot 0.200 0.085 0.142 [0.130, 0.154] <0.001 25%
Claude-3-5-Sonnet-20240620 Anthropic News Scratchpad 0.163 0.127 0.145 [0.134, 0.156] <0.001 19%
Imputed Forecaster ForecastBench 0.250 0.045 0.147 [0.140, 0.154] <0.001 24%
GPT-4 OpenAI Freeze values Zero shot 0.212 0.083 0.147 [0.136, 0.159] <0.001 23%
Gemini-1.5-Pro Google Freeze values Zero shot 0.207 0.090 0.148 [0.136, 0.161] <0.001 21%
GPT-4 OpenAI Freeze values Scratchpad 0.191 0.106 0.149 [0.137, 0.161] <0.001 19%
Claude-3-Opus-20240229 Anthropic Freeze values Scratchpad 0.187 0.113 0.150 [0.139, 0.161] <0.001 18%
Claude-3-5-Sonnet-20240620 Anthropic News Superforecaster 2 0.172 0.129 0.150 [0.138, 0.162] <0.001 17%
LLM Crowd ForecastBench News 0.238 0.064 0.151 [0.144, 0.158] <0.001 19%
GPT-4-Turbo-2024-04-09 OpenAI News with freeze values Scratchpad 0.210 0.095 0.152 [0.141, 0.163] <0.001 22%
LLM Crowd ForecastBench News 0.239 0.067 0.153 [0.146, 0.160] <0.001 19%
LLM Crowd ForecastBench News 0.240 0.066 0.153 [0.146, 0.160] <0.001 19%
Mistral-Large-Latest Mistral AI Freeze values Scratchpad 0.186 0.121 0.154 [0.144, 0.164] <0.001 16%
Llama-3-70b-Chat-Hf Meta Freeze values Zero shot 0.196 0.112 0.154 [0.141, 0.167] <0.001 20%
Gemini-1.5-Flash Google Freeze values Scratchpad 0.178 0.132 0.155 [0.142, 0.168] <0.001 19%
Gemini-1.5-Pro Google News Superforecaster 1 0.178 0.132 0.155 [0.144, 0.166] <0.001 19%
Llama-3-70b-Chat-Hf Meta Freeze values Scratchpad 0.203 0.108 0.156 [0.146, 0.166] <0.001 17%
Qwen1.5-110B-Chat Qwen Scratchpad 0.172 0.140 0.156 [0.147, 0.165] <0.001 15%
Gemini-1.5-Pro Google Zero shot 0.207 0.106 0.157 [0.145, 0.169] <0.001 18%
Gemini-1.5-Flash Google Scratchpad 0.178 0.136 0.157 [0.146, 0.168] <0.001 15%
GPT-4 OpenAI Scratchpad 0.191 0.123 0.157 [0.148, 0.166] <0.001 15%
GPT-4-Turbo-2024-04-09 OpenAI Zero shot 0.200 0.117 0.159 [0.147, 0.170] <0.001 19%
Claude-2.1 Anthropic Scratchpad 0.227 0.091 0.159 [0.149, 0.168] <0.001 20%
Gemini-1.5-Flash Google Freeze values Zero shot 0.217 0.103 0.160 [0.146, 0.174] <0.001 23%
GPT-4-Turbo-2024-04-09 OpenAI News Scratchpad 0.210 0.109 0.160 [0.149, 0.170] <0.001 17%
Claude-3-Opus-20240229 Anthropic Zero shot 0.178 0.144 0.161 [0.148, 0.175] <0.001 19%
GPT-4-Turbo-2024-04-09 OpenAI News Superforecaster 2 0.204 0.120 0.162 [0.151, 0.173] <0.001 18%
Qwen1.5-110B-Chat Qwen News with freeze values Scratchpad 0.199 0.125 0.162 [0.152, 0.172] <0.001 16%
Mistral-Large-Latest Mistral AI Freeze values Zero shot 0.201 0.125 0.163 [0.149, 0.176] <0.001 19%
Claude-3-5-Sonnet-20240620 Anthropic Zero shot 0.177 0.151 0.164 [0.150, 0.178] <0.001 18%
GPT-4o OpenAI News Superforecaster 3 0.206 0.125 0.166 [0.156, 0.176] <0.001 17%
Claude-2.1 Anthropic Freeze values Scratchpad 0.227 0.105 0.166 [0.154, 0.177] <0.001 21%
Notes:
1. This shows performance on all 200 standard questions from the human question set plus those combination questions from the LLM question set where humans
provided forecasts on both components ($Q1$ and $Q2$). LLM scores are only for this combined question set. Human forecasts for combination questions are generated
from their forecasts on the component questions by assuming independence (which is not always the case, putting humans at a disadvantage).
Appendix I Human Survey Demographics
We collected demographic information from the 500 human forecasters in the public survey. Summaries of participants’ age, gender, ethnicity, and country of residence are shown in the tables below.
Table 21: Age Distribution
Age Percentage
18-24 years old 32.0%
25-34 years old 43.4%
35-44 years old 14.4%
45-54 years old 5.4%
Over 55 4.8%
Table 22: Gender Distribution
Gender Percentage
Male 53.4%
Female 46.2%
Prefer not to say 0.4%
Table 23: Ethnicity Distribution
Ethnicity Percentage
White 48.6%
Black 33.4%
Mixed 8.6%
Asian 4.4%
Other 3.4%
Prefer not to say 1.6%
Table 24: Country of Residence Distribution
Country Percentage
South Africa 31.2%
United States 15.4%
Poland 9.0%
Portugal 7.8%
United Kingdom 4.6%
Mexico 4.2%
Chile 3.8%
Italy 3.4%
Greece 3.0%
Hungary 3.0%
(25 others) 14.6%
We did not collect similar demographic information from the superforecasters participating in the study, but are reasonably certain that the superforecasters in this study are roughly representative of superforecasters as a whole. Describing forecasters previously recruited by Good Judgment Project, Mellers et al. (2015) noted that participants "tended to be men (83%) and U.S. citizens (74%), with an average age of 40 years."
Appendix J Performance Breakdown
In tables Table 25 and Table 26 we look into how the best-performing LLM, Claude 3.5 Sonnet (Scratchpad prompt with freeze values), and Superforecasters performed when assessing performance by category and by forecast horizon.
In Table 25, we see that Claude 3.5 Sonnet slightly Superforecasters on Environment & Energy questions and Sports questions.
As a reminder, we ask for forecasts on dataset questions at 8 diffreent forecast horizons, as explained in Section 3.2. Here, we breakdown performance at the forecast horizons that have resolved by the publication date: 7 days, 30 days, and 90 days. We see that Superforecasters outperformed Claude 3.5 Sonnet at all forecast horizons.
Table 25: Brier score by category
Category N Claude 3.5 Sonnet Superforecaster Difference
Arts & Recreation 8 0.086 0.053 0.033
Economics & Business 137 0.188 0.105 0.083
Environment & Energy 63 0.145 0.158 -0.013
Healthcare & Biology 37 0.003 0.001 0.002
Politics & Governance 17 0.061 0.017 0.044
Science & Tech 15 0.160 0.139 0.021
Security & Defense 73 0.049 0.027 0.022
Sports 43 0.020 0.034 -0.014
Table 26: Brier score by forecast horizon
Forecast Horizon N Claude 3.5 Sonnet Superforecaster Difference
7-day 104 0.159 0.132 0.027
30-day 106 0.150 0.138 0.012
90-day 183 0.098 0.083 0.015